# Sample Compression Unleashed: New Generalization Bounds for Real Valued Losses

### v2

@@ 3 A General Sample-Compress Bound
For any distribution $\operatorname{\mathcal{D}}$ over $\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}$, for any family of set of messages $\{M(\operatorname{\mathbf{i}})\,|\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)\}$, for any deterministic reconstruction function $\operatorname{\mathscr{R}}$ that outputs sample-compressed predictors $h\in\operatorname{\mathcal{H}}$, for any loss :×$\ell:\operatorname{\mathcal{H}}\times\operatorname{\mathcal{X}}\times Y\to\operatorname{\mathbb{R}}$, for any comparator function $\Delta:\operatorname{\mathbb{R}}\times\operatorname{\mathbb{R}}\to\operatorname{\mathbb{R}}$ and for any $\delta\in(0,1]$, with probability at least $1-\delta$ over the draw of $S\sim\operatorname{\mathcal{D}}^{n}$, we have
@@
$~{}~{}~{}~{}\leq\frac{1}{|\operatorname{\mathbf{i}}^{c}|}\quantity[\log\matrixquantity(n\\ \operatorname{|\mathbf{i}|})+\log\quantity(\frac{\mathcal{E}_{\Delta}\quantity(\mathbf{i},\sigma)}{\zeta(\operatorname{|\mathbf{i}|})P_{M(\operatorname{\mathbf{i}})}(\sigma)\delta})]$
@@
$=e^{-t\epsilon}\operatorname*{\mathbb{E}}_{S_{\operatorname{\mathbf{i}}}\sim\operatorname{\mathcal{D}}^{n}}\operatorname*{\mathbb{E}}_{S_{\operatorname{\mathbf{i}}^{c}}\sim\operatorname{\mathcal{D}}^{n}}e^{t\Delta\quantity(\operatorname{\widehat{\mathcal{L}}}_{S_{\operatorname{\mathbf{i}}^{c}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{i}}},\sigma)),\mathcal{L}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{i}}},\sigma)))}$
@@
$\delta_{\operatorname{\mathbf{i}}}^{\sigma}=e^{-t\epsilon}\operatorname*{\mathbb{E}}_{S_{\operatorname{\mathbf{i}}}\sim\operatorname{\mathcal{D}}^{n}}\operatorname*{\mathbb{E}}_{S_{\operatorname{\mathbf{i}}^{c}}\sim\operatorname{\mathcal{D}}^{n}}e^{t\Delta\quantity(\operatorname{\widehat{\mathcal{L}}}_{S_{\operatorname{\mathbf{i}}^{c}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{i}}},\sigma)),\mathcal{L}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{i}}},\sigma)))}$ (2)
and solve for $\epsilon$, using $t=n-\operatorname{|\mathbf{i}|}$. The obtained solution is used to replace the $\epsilon$ in Eq. 1, which gives a bound valid with probability $\delta_{\operatorname{\mathbf{i}}}^{\sigma}$ for every single predictor $\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{i}}},\sigma)$. By setting $\delta_{\operatorname{\mathbf{i}}}^{\sigma}=P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{M(\operatorname{\mathbf{i}})}(\sigma)\delta$ and applying a union bound over all $\mathbf{i}\in\operatorname{\mathscr{P}}(n)$, $\sigma\in M(\operatorname{\mathbf{i}})$, the final result holds uniformly with probability $\delta$ for all predictors outputted by $\operatorname{\mathscr{R}}$. ∎
Theorem 3 holds for any comparator function $\Delta$ such that $\mathcal{E}_{\Delta}$ is finite for any pair $(\operatorname{\mathbf{i}},\sigma)$. Although bounding $\operatorname{\mathcal{E}}_{\Delta}$ can be challenging, it was extensively studied for convex functions in PAC-Bayesian theory (e.g., McAllester,, 1998; Maurer,, 2004; Casado et al.,, 2024; Hellström and Guedj,, 2024). We leverage this theory and present novel corollaries for the three most well-known comparators.
@@
In the setting of Theorem 3, for any $C>0$, for a loss function :×$\ell:\operatorname{\mathcal{H}}\times\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}\to[0,1]$, with probability at least $1-\delta$ over the draw of $S\sim\operatorname{\mathcal{D}}^{n}$, we have
$\forall\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n),\sigma\in M(\operatorname{\mathbf{i}}):\,\mathcal{L}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{i}}},\sigma))\leq\frac{1-\exp(-\epsilon_{C}(\operatorname{\mathbf{i}},\sigma,\delta))}{1-e^{-C}}$
@@
For $0\leq q,p\leq 1$, there exists $C^{*}=\operatorname*{arg\,sup}_{C>0}\Delta_{C}(q,p)$ such that $\Delta_{C^{*}}$ gives the tightest PAC-Bayesian bounds (Foong et al.,, 2021). This result also holds true for Theorem 3, when restricted to proper, convex and lower semicontinuous comparator functions $\Delta:[0,1]\times[0,1]\to\operatorname{\mathbb{R}}$. Unfortunately, the hyperparameterized bounds hold for only one value of $C$, chosen prior to seeing $S$. With a union bound argument, we can consider multiple parameters $C$ simultaneously, but there is no guarantee that $C^{*}$ is in this set. To circumvent this problem, we can use the binary Kullback-Leibler divergence comparator function $\mathrm{kl}(q,p)=q\ln\frac{q}{p}+(1-q)\ln\frac{1-q}{1-p},$ which is equivalent to $\Delta_{C^{*}}$, as per the following proposition.
@@
In practice, even with the term $1=\operatorname{\mathcal{E}}_{\Delta_{C}}(n)\leq\operatorname{\mathcal{E}}_{\mathrm{kl}}(n)=2\sqrt{n}$, the $\mathrm{kl}$ bound will usually yield tighter bounds than Corollary 4, as the optimal value of $C$ is unlikely to be selected before computing the bound. Moreover, the $\mathrm{kl}$ is known to be optimal for $[0,1]$-valued losses, as per the results of Hellström and Guedj, (2024).
@@
In the setting of Theorem 3, for a loss function :×$\ell:\operatorname{\mathcal{H}}\times\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}\to[0,1]$, with probability at least $1-\delta$ over the draw of $S\sim\operatorname{\mathcal{D}}^{n}$, we have
@@
$\mathcal{L}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{i}}},\sigma))\leq\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S_{\operatorname{\mathbf{i}}^{c}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{i}}},\sigma)),\epsilon_{\mathrm{kl}}(\operatorname{\mathbf{i}},\sigma,\delta))$
@@
This result could be extended to the hypothesis-dependent range condition of Haddouche et al., (2021), any unbounded losses under model-dependent assumptions (Casado et al.,, 2024) or more general tail behaviors (Rodríguez-Gálvez et al.,, 2024). Note that this result encompasses bounded losses with a range of $[a,b]$, as they are sub-Gaussian with $\varsigma=\frac{b-a}{2}$.
@@ 3.1 Behavior in the consistent case
In this section, we present a new theoretical results that justifies the tightness of the bounds observed in Section 4. In this setting, predictors stop training when the empirical error reaches zero. Indeed, we show that when $\widehat{R}_{S_{\operatorname{\mathbf{i}}^{c}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{i}}},\sigma))=0$, our new bounds from Corollaries 4 and 6 are extremely tight upper bounds to the binomial bound of Laviolette et al., (2005).
@@
In the consistent case, i.e. when the training error is zero, Corollary 4 is arbitrarily close to the binomial tail inversion of Theorem 1. Moreover, Corollary 6 is a tight upper bound up to a constant $K(m,\delta)$ that decreases for $m$ large enough and tends to 00 when $m$ tends to $\infty$. Indeed,
$\overline{\rm{Bin}}\quantity(0,m,\delta)\ =$ $1-\exp\quantity(\frac{-1}{m}\ln\frac{1}{\delta})$ (3)
$=$ $\lim_{C\to\infty}\frac{1}{1-e^{-C}}\quantity[1-\exp\quantity(-\frac{1}{m}\ln\frac{1}{\delta})]$ (4)
$=$ $\inf_{C>0}\frac{1}{1-e^{-C}}\quantity[1-\exp\quantity(-\frac{1}{m}\ln\frac{1}{\delta})]$ (5)
$=$ $\mathrm{kl}^{-1}\quantity(0,\frac{1}{m}\ln\frac{1}{\delta})$ (6)
$\leq$ $\mathrm{kl}^{-1}\quantity(0,\frac{1}{m}\ln\frac{2\sqrt{m}}{\delta}).$ (7)
We prove the previous sequence of results in Section C.3. In the previous result, we relate the analytical form of $\overline{\textrm{Bin}}\quantity(0,m,\delta)$ in Eq. 3 to the bound of Corollary 4. This relation is known to hold as an inequality for any $k\geq 0$, but we show that it is an equality for $k=0$. In Eq. 4, we show that the hyperparameterized distance is arbitrarily tight to the binomial tail inversion, as it approaches $\overline{\textrm{Bin}}\quantity(0,m,\delta)$ when $C\to\infty$. We show in Eq. 5 that the minimal value of the bound is obtained when $C$ tends to $\infty$. In Eq. 6, we relate the hyperparameterized distance to the Kullback-Leibler divergence, which we then upper bound in Eq. 7, achieving the result of Corollary 6. Finally, in Eq. 8, we demonstrate that Corollary 6 is a tight upper bound to the binomial tail inversion, up to a constant $K(m,\delta)$ that decreases when $m$ is large and tends to 00 when $n\to\infty$. For example, with $\delta=0.01$, $K(m,\delta)$ is bounded by $K(7,0.01)\approx 0.11$, decreases for $n\geq 8$ and reaches $K(m,\delta)\leq 0.01$ at $m=357$.

### v3

@@ 3 A GENERAL SAMPLE-COMPRESSION BOUND
For any distribution $\operatorname{\mathcal{D}}$ over $\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}$, for any family of set of messages $\{M(\operatorname{\mathbf{i}})\,|\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)\}$, for any deterministic reconstruction function $\operatorname{\mathscr{R}}$ that outputs sample-compressed predictors $h\in\overline{\operatorname{\mathcal{H}}}$, for any loss :×$\ell:\operatorname{\mathcal{H}}\times\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}\to\operatorname{\mathbb{R}}$, for any comparator function $\Delta:\operatorname{\mathbb{R}}\times\operatorname{\mathbb{R}}\to\operatorname{\mathbb{R}}$ and for any $\delta\in(0,1]$, with probability at least $1-\delta$ over the draw of $S\sim\operatorname{\mathcal{D}}^{n}$, we have
@@
$~{}~{}~{}~{}\leq\frac{1}{|\operatorname{\mathbf{i}}^{c}|}\quantity[\log\matrixquantity(n\\ \operatorname{|\mathbf{i}|})+\log\quantity(\frac{\mathcal{E}_{\Delta}\quantity(\mathbf{i},\sigma)}{\zeta(\operatorname{|\mathbf{i}|})P_{M(\operatorname{\mathbf{i}})}(\sigma)\delta})]\,,$
@@
$=e^{-t\epsilon}\operatorname*{\mathbb{E}}_{S_{\operatorname{\mathbf{i}}}\sim\operatorname{\mathcal{D}}^{\operatorname{|\mathbf{i}|}}}\operatorname*{\mathbb{E}}_{S_{\operatorname{\mathbf{i}}^{c}}\sim\operatorname{\mathcal{D}}^{|\mathbf{i}^{c}|}}e^{t\Delta\quantity(\operatorname{\widehat{\mathcal{L}}}_{S_{\operatorname{\mathbf{i}}^{c}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{i}}},\sigma)),\mathcal{L}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{i}}},\sigma)))}$
@@
$\delta_{\operatorname{\mathbf{i}}}^{\sigma}=e^{-t\epsilon}\operatorname*{\mathbb{E}}_{S_{\operatorname{\mathbf{i}}}\sim\operatorname{\mathcal{D}}^{\operatorname{|\mathbf{i}|}}}\operatorname*{\mathbb{E}}_{S_{\operatorname{\mathbf{i}}^{c}}\sim\operatorname{\mathcal{D}}^{|\mathbf{i}^{c}|}}e^{t\Delta\quantity(\operatorname{\widehat{\mathcal{L}}}_{S_{\operatorname{\mathbf{i}}^{c}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{i}}},\sigma)),\mathcal{L}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{i}}},\sigma)))}$ (2)
and solve for $\epsilon$, using $t=|\mathbf{i}^{c}|$. The obtained solution is used to replace the $\epsilon$ in Eq. 1, which gives a bound valid with probability $\delta_{\operatorname{\mathbf{i}}}^{\sigma}$ for every single predictor $\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{i}}},\sigma)$. By setting $\delta_{\operatorname{\mathbf{i}}}^{\sigma}=P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{M(\operatorname{\mathbf{i}})}(\sigma)\delta$ and applying a union bound over all $\mathbf{i}\in\operatorname{\mathscr{P}}(n)$, $\sigma\in M(\operatorname{\mathbf{i}})$, the final result holds uniformly with probability $\delta$ for all predictors outputted by $\operatorname{\mathscr{R}}$. ∎
Theorem 3 holds for any comparator function $\Delta$ such that $\mathcal{E}_{\Delta}$ is finite for any pair $(\operatorname{\mathbf{i}},\sigma)$. Although bounding $\operatorname{\mathcal{E}}_{\Delta}$ can be challenging, it was extensively studied for convex functions in PAC-Bayesian theory (e.g., McAllester,, 1998; Maurer,, 2004; Casado Telletxea et al.,, 2025; Hellström and Guedj,, 2024). We leverage this theory and present novel corollaries for the three most well-known comparators.
@@
In the setting of Theorem 3, for any $C>0$, for any loss function :×$\ell:\operatorname{\mathcal{H}}\times\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}\to[0,1]$, with probability at least $1-\delta$ over the draw of $S\sim\operatorname{\mathcal{D}}^{n}$, we have
$\forall\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n),\sigma\in M(\operatorname{\mathbf{i}}):\,\mathcal{L}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{i}}},\sigma))\leq\frac{1-\exp(-\epsilon_{C}(\operatorname{\mathbf{i}},\sigma,\delta))}{1-e^{-C}},$
@@
For $0\leq q,p\leq 1$, there exists $C^{*}=\operatorname*{arg\,sup}_{C>0}\Delta_{C}(q,p)$ such that $\Delta_{C^{*}}$ gives the tightest PAC-Bayesian bounds (Foong et al.,, 2021). This result also holds true for Theorem 3, when restricted to proper, convex and lower semicontinuous comparator functions $\Delta:[0,1]\times[0,1]\to\operatorname{\mathbb{R}}$. Unfortunately, the $\Delta_{C}$ bound hold for only one value of $C$, chosen prior to seeing $S$. With a union bound argument, we can consider multiple parameters $C$ simultaneously, but there is no guarantee that $C^{*}$ is in this set. To circumvent this problem, we can use the binary Kullback-Leibler divergence comparator function $\mathrm{kl}(q,p)=q\ln\frac{q}{p}+(1-q)\ln\frac{1-q}{1-p},$ which is equivalent to $\Delta_{C^{*}}(q,p)$, as per the following proposition.
@@
In practice, even with the term $1=\operatorname{\mathcal{E}}_{\Delta_{C}}(\operatorname{\mathbf{i}},\sigma)\leq\operatorname{\mathcal{E}}_{\mathrm{kl}}(\operatorname{\mathbf{i}},\sigma)=2\sqrt{n-\operatorname{|\mathbf{i}|}}$, the $\mathrm{kl}$ bound stated below (Corollary 6) usually yield tighter bounds than the $\Delta_{C}$ bound (Corollary 4), as the optimal value $C^{*}$ is unlikely to be selected before computing the bound. Moreover, the $\mathrm{kl}$ is known to be optimal for $[0,1]$-valued losses, as per the results of Hellström and Guedj, (2024).
@@
In the setting of Theorem 3, for any loss function :×$\ell:\operatorname{\mathcal{H}}\times\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}\to[0,1]$, with probability at least $1-\delta$ over the draw of $S\sim\operatorname{\mathcal{D}}^{n}$, we have
@@
$\mathcal{L}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{i}}},\sigma))\leq\mathrm{kl}^{-1}\quantity(\operatorname{\widehat{\mathcal{L}}}_{S_{\operatorname{\mathbf{i}}^{c}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{i}}},\sigma)),\epsilon_{\mathrm{kl}}(\operatorname{\mathbf{i}},\sigma,\delta)),$
@@
Note that this result encompasses bounded losses with a range of $[a,b]$, as they are sub-Gaussian with $\varsigma=\frac{b-a}{2}$. It can be extended to the hypothesis-dependent range condition of Haddouche et al., (2021), any unbounded losses under model-dependent assumptions (Casado Telletxea et al.,, 2025) or more general tail behaviors (Rodríguez-Gálvez et al.,, 2024).
Table 1: Results for the CNNs trained using P2L on the binary MNIST problems. The results displayed obtained the tightest P2L bound. All metrics presented are in percent (%).
Dataset Validation error Test error $\mathrm{kl}$ bound Binomial bound P2L bound $\operatorname{|\mathbf{i}|}/n$ Baseline test error
MNIST08 0.33$\pm$0.17 0.25$\pm$0.10 5.05$\pm$0.16 5.00$\pm$0.16 1.04$\pm$0.04 0.87$\pm$0.03 0.22$\pm$0.05
MNIST17 0.20$\pm$0.08 0.38$\pm$0.16 4.33$\pm$0.21 4.29$\pm$0.21 0.86$\pm$0.05 0.72$\pm$0.04 0.17$\pm$0.08
MNIST23 0.39$\pm$0.12 0.27$\pm$0.10 8.20$\pm$0.34 8.15$\pm$0.34 1.86$\pm$0.09 1.61$\pm$0.09 0.16$\pm$0.05
MNIST49 0.82$\pm$0.11 0.77$\pm$0.17 10.52$\pm$0.37 10.47$\pm$0.37 2.53$\pm$0.11 2.23$\pm$0.10 0.44$\pm$0.08
MNIST56 0.46$\pm$0.12 0.47$\pm$0.15 6.29$\pm$0.22 6.24$\pm$0.22 1.35$\pm$0.06 1.15$\pm$0.05 0.30$\pm$0.05
Table 2: Results for the CNNs trained using P2L on the binary MNIST problems and stopped at the iteration with the minimum $\mathrm{kl}$ bound. The results displayed obtained the tightest $\mathrm{kl}$ bound. Metrics are in percents (%).
Dataset Validation error Test error $\mathrm{kl}$ bound Binomial bound Train error $\operatorname{|\mathbf{i}|}/n$ Baseline test error
MNIST08 0.49$\pm$0.39 0.49$\pm$0.26 4.71$\pm$0.25 5.33$\pm$0.62 0.24$\pm$0.23 0.62$\pm$0.14 0.22$\pm$0.05
MNIST17 0.45$\pm$0.18 0.48$\pm$0.11 3.70$\pm$0.21 4.37$\pm$0.11 0.23$\pm$0.08 0.43$\pm$0.08 0.17$\pm$0.08
MNIST23 0.74$\pm$0.28 0.84$\pm$0.21 6.56$\pm$0.38 8.09$\pm$0.64 0.64$\pm$0.32 0.77$\pm$0.20 0.16$\pm$0.05
MNIST49 1.16$\pm$0.31 1.13$\pm$0.24 8.60$\pm$0.46 9.61$\pm$0.68 0.51$\pm$0.28 1.26$\pm$0.23 0.44$\pm$0.08
MNIST56 0.94$\pm$0.09 0.70$\pm$0.20 5.42$\pm$0.31 6.49$\pm$0.81 0.43$\pm$0.23 0.65$\pm$0.10 0.30$\pm$0.05
@@ 3.1 Behavior in the consistent case
In this section, we present new theoretical results that justify the tightness of the bounds observed in Section 4, in which we train different types of models with P2L. By construction, P2L is designed to stop when the complement error $\widehat{R}_{S_{\operatorname{\mathbf{i}}^{c}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{i}}},\sigma))$ is zero. In this consistent setting, where the predictor always finishes training with zero errors, we can prove that Corollaries 4 and 6 are tight upper bounds of Theorem 1.
The first result, presented in Theorem 8, states that the $\Delta_{C}$ bound is an arbitrarily tight upper bound of the binomial tail inversion bound of Laviolette et al., (2005). Indeed, in the following theorem, we show that Corollary 4 is minimized by $C$ tending to $\infty$ and is equal to Theorem 1 in the limit of $C\to\infty$.
@@
In the consistent case, i.e. when $\widehat{R}_{S_{\operatorname{\mathbf{i}}^{c}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{i}}},\sigma))=0$, Corollary 4 is arbitrarily close to the binomial tail inversion of Theorem 1. Indeed, we have
$\overline{\rm{Bin}}\quantity(0,|\operatorname{\mathbf{i}}^{c}|,\delta_{\operatorname{\mathbf{i}}}^{\sigma})$ $=\inf_{C>0}\left\{\frac{1-\exp\quantity(-\frac{1}{|\operatorname{\mathbf{i}}^{c}|}\ln\frac{1}{\delta_{\operatorname{\mathbf{i}}}^{\sigma}})}{1-e^{-C}}\right\}$ (3)
$=\lim_{C\to\infty}\left\{\frac{1-\exp\quantity(-\frac{1}{|\operatorname{\mathbf{i}}^{c}|}\ln\frac{1}{\delta_{\operatorname{\mathbf{i}}}^{\sigma}})}{1-e^{-C}}\right\}$ (4)
with $\delta_{\operatorname{\mathbf{i}}}^{\sigma}=\smallmatrixquantity(n\\ \operatorname{|\mathbf{i}|})^{-1}\zeta(\operatorname{|\mathbf{i}|})P_{M(\operatorname{\mathbf{i}})}(\sigma)\delta$.
The following theorem states that the Kullback-Leibler divergence bound is a tight upper bound of the binomial tail inversion bound, up to a constant.
Theorem 9.
In the consistent case, i.e. when $\widehat{R}_{S_{\operatorname{\mathbf{i}}^{c}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{i}}},\sigma))=0$, Corollary 6 is a tight upper bound of Theorem 1 up to a constant $K(m,\delta)$. Indeed, we have
$\overline{\rm{Bin}}\quantity(0,|\operatorname{\mathbf{i}}^{c}|,\delta_{\operatorname{\mathbf{i}}}^{\sigma})\ \leq$ $\ \mathrm{kl}^{-1}\quantity(0,\frac{1}{|\operatorname{\mathbf{i}}^{c}|}\ln\frac{2\sqrt{|\operatorname{\mathbf{i}}^{c}|}}{\delta_{\operatorname{\mathbf{i}}}^{\sigma}})$ (5)
$=$ $\ \overline{\rm{Bin}}\quantity(0,|\operatorname{\mathbf{i}}^{c}|,\delta_{\operatorname{\mathbf{i}}}^{\sigma})+K(|\operatorname{\mathbf{i}}^{c}|,\delta_{\operatorname{\mathbf{i}}}^{\sigma})\,,$ (6)
with $K(m,\delta)\ =\ \exp\quantity(-\tfrac{1}{m}\ln\tfrac{1}{\delta})-\exp\quantity(-\tfrac{1}{m}\ln\tfrac{2\sqrt{m}}{\delta})$ and $\delta_{\operatorname{\mathbf{i}}}^{\sigma}=\smallmatrixquantity(n\\ \operatorname{|\mathbf{i}|})^{-1}\zeta(\operatorname{|\mathbf{i}|})P_{M(\operatorname{\mathbf{i}})}(\sigma)\delta$.
The proofs of both Theorems 8 and 9 can be found in Section C.3.
Note that the constant $K(m,\delta)$ of Equation (6) tends to 00 when $m$ tends to $\infty$ and is bounded by
$0\ \leq\ K(m,\delta)\ \leq\ \frac{1}{m}\ln\frac{2\sqrt{m}}{\delta}.$
With $\delta=0.01$, $K(m,\delta)$ is maximized at $m\approx 7.35$, with $K(7.35,0.01)\approx 0.11$. However, as depicted by the empirical results of Section 4, $\delta_{\operatorname{\mathbf{i}}}^{\sigma}$ is orders of magnitude smaller than $0.01$ in practical situations. [^2] Given a dataset of $10597$ datapoints, a compression set of size $92$ and $\delta=0.01$, we have $K(10505,10^{-234})\approx 0.0005$. This is indeed the difference that we observe in Table 1 between the $\mathrm{kl}$ bound and the binomial bound for MNIST08.